The length and width of the rectangular bottom of Sabrina's new fish tank is twice the length and width of the rectangular bottom of her old tank. The area of the bottom of the new tank is 16 ft^2. The lengh of the old tank was 3 feet more than the width of the new tank.
a. Using F to represent the width of the old tank, write an equation to show the area of the new tank bottom
b. What was the width of the old tank?
c. What is the length of the new tank?

Answer by Alan3354(69443)

The length and width of the rectangular bottom of Sabrina's new fish tank is twice the length and width of the rectangular bottom of her old tank. The area of the bottom of the new tank is 16 ft^2.
The lengh of the old tank was 3 feet more than the width of the new tank.
=======
-----
The area of the old tank is 16/4 = 4 sq ft.
---
The old tank bottom was 1 by 4 feet.
---
a. Using F to represent the width of the old tank, write an equation to show the area of the new tank bottom
Area = F*2*(L*2) = 16
----
b. What was the width of the old tank?
1 foot
------
c. What is the length of the new tank?
8 ft
==================
A poorly written problem, btw.